OpenStudy (alyssajobug):

Multiple hard Algebra 2 questions ------------------------------------------------------------------- The pilot of a helicopter plans to release a bucket of water on a forest fire. The height y in feet of the water t seconds after its release is modeled by y = -16t^2 - 2t + 400. The horizontal distance x in feet between the water and its point of release is modeled by x = 91t. To the nearest foot, at what horizontal distance from the target should the pilot begin releasing the water? in feet -------------------------------------------------------------------- If a tightrope walker falls, he will land on a safety net. His height h in feet after a fall can be modeled by h(t) = 60 - 16t2, where t is the time in seconds. If the safety net is 11 feet above the ground, how many seconds will the tightrope walker fall before landing on the safety net? Round your answer to the nearest hundredth. in seconds

OpenStudy (phi):

do they tell you the height of the helicopter ?

OpenStudy (alyssajobug):

no, this is all they give me :-/

OpenStudy (phi):

I guess they want you to figure out "t" when y=0 y=0 means the water is at height 0. you want to know how long it takes. once you find t, use it in x = 91*t to find x so step 1 is solve for t in -16t^2 - 2t + 400=0

OpenStudy (alyssajobug):

so do I use the quadratic equasion then? @phi

OpenStudy (phi):

that looks like your best approach. I would first divide the equation (all terms, both sides) by 2, to make it just a little simpler. you get -8 t^2 -t + 200 = 0/2 or \[ -8 t^2 -t + 200 = 0 \] a= -8, b= -1, c =+200

OpenStudy (alyssajobug):

\[x=1\pm \frac{ \sqrt{-1^{2}-4\times-8\times200}\ }{ 2\times-8 }\]

OpenStudy (phi):

ok, except the 2*-8 divides the 1 also.

OpenStudy (alyssajobug):

\[x=1\pm \frac{-6399 }{ -16 }\]

OpenStudy (alyssajobug):

what do uou mean "ok, except the 2*-8 divides the 1 also."?

OpenStudy (alyssajobug):

\[x =-b \pm \frac{ \sqrt{b ^{2-4\times a \times c}} }{ 2a }\]

OpenStudy (phi):

\[ x= \frac{ 1\pm\sqrt{(-1)^{2}-4\times-8\times200}\ }{ 2\times-8 } \]

OpenStudy (phi):

see http://www.purplemath.com/modules/solvquad4.htm

OpenStudy (phi):

now inside the square root, you should get -1*-1 + 4*8*200

OpenStudy (alyssajobug):

oh, ok so from there it is \[x =1\pm \frac{ \sqrt{1+6,400} }{ -16 }\]

OpenStudy (alyssajobug):

\[=1\pm \frac{ \sqrt{6,401} }{ -16 }\]

OpenStudy (phi):

you are a bit stubborn, aren't you ?? it is \[ = \frac{ 1\pm\sqrt{6,401} }{ -16 } \]

OpenStudy (alyssajobug):

that cant be the final form, where does it go from there? where do I even begin to simplify it from there?

OpenStudy (phi):

This one needs a calculator.

OpenStudy (alyssajobug):

ok

OpenStudy (phi):

only 1 of the solutions (the positive one) makes sense.

OpenStudy (alyssajobug):

ok so \[\sqrt{6401} =80.0062497559\] so it would now read \[=\frac{ 1\pm80.0062497559 }{ -16 }\]

OpenStudy (phi):

notice, to get a positive time, you want the top to be negative. in other words choose (1-80.0062497559)/-16

OpenStudy (alyssajobug):

oh! ok, so the solutions are -5.063125 and 4.938125 so the 4.9 would be the best answer?

OpenStudy (phi):

I would not round the time. only round after you get the final number.

OpenStudy (alyssajobug):

oh ok so just input the whole 4.938125 then?

OpenStudy (phi):

now find x x = 91t

OpenStudy (alyssajobug):

x = 449.369375

OpenStudy (alyssajobug):

now do I round it? to 449.4

OpenStudy (phi):

To the nearest foot, at what horizontal distance

OpenStudy (alyssajobug):

oh ok so just to 449 then, ok

OpenStudy (phi):

btw, you got t= 4.938125 however, the correct t is 4.9378906 Your answer was close enough, but I would re-do that step... just to make sure you can get the correct answer.

OpenStudy (alyssajobug):

ok, my lesson said it was right the 449, ya, my iphone only calculates to the 7th place sadly, I need to get a real calculator eventually

OpenStudy (alyssajobug):

it rounds what is after it

OpenStudy (alyssajobug):

oh, not sure what it did there, thats the same amount of digits, I'll have to watch that with this thing

OpenStudy (phi):

7 places is fine.... I just wanted to make sure you weren't doing something wrong. As for the second problem, half of the "fun" is figuring out what equation to solve. Can you try to write down the equation you should solve?

OpenStudy (alyssajobug):

ya, I think so, it would be \[-16t ^{2}+60\] right?

OpenStudy (phi):

that is not an equation. equations have an = sign in them

OpenStudy (alyssajobug):

well, \[h(t)=-16t ^{2}+60\] then?

OpenStudy (alyssajobug):

no, that equasion =0 then?

OpenStudy (phi):

that tells you His height h in feet after a fall, after t seconds what is his height when he hits the net ?

OpenStudy (phi):

−16t^2+60=0 that would let you find the time it takes to hit the ground (the ground is at "height" zero)

OpenStudy (phi):

First, we do not (yet) have the correct equation. Second, you are *STILL* using the wrong formula for the quadratic equation.

OpenStudy (phi):

what is his height when he hits the net ?

OpenStudy (alyssajobug):

ya, the height of the net is 11 ft off the ground so does that mean that it is -11x in the equasion?

OpenStudy (alyssajobug):

so the equasion reads -16t^2-11x+60=0?

OpenStudy (phi):

This is important... h(t) = -16t^2 + 60 that tells you his height after t seconds. for example, when t=0 (just when he starts to fall) his height is h(0)= -16*0+60 or 60 feet. we know have some number of seconds his height will be less (he is falling) in other words we know that h(t) = 11 feet we want to find the time when that happens.

OpenStudy (phi):

we know after some number of seconds his height will be less (he is falling) in other words we know that at some point h(t) = 11 feet we want to find the time when that happens.

OpenStudy (alyssajobug):

ok so would 11 replace the 0 then? so -16t^2+60=11?

OpenStudy (phi):

yes, exactly. I hope that makes sense.

OpenStudy (alyssajobug):

ya, I think it does, so to make it quadratic able, I need to make it =0 so I take 11 from both sides?

OpenStudy (alyssajobug):

-16t^2+49=0?

OpenStudy (phi):

yes

OpenStudy (alyssajobug):

ok so put that into quadratic equasion is \[x=\frac{ 0\pm \sqrt{b ^{2}-4\times-16\times49} }{ 2\times-16 }\]?

OpenStudy (alyssajobug):

whoops should have been a 0 in that b, was too busy singin the song in my head!

OpenStudy (phi):

that will work. fyi, notice when you don't have the "x" term, you can do this \[-16t^2+49=0 \\ -16t^2 = -49 \\ t^2= \frac{49}{16} \]

OpenStudy (phi):

which can be solved in your head (if the song is not confusing you)

OpenStudy (alyssajobug):

t^2=2.0625

OpenStudy (alyssajobug):

no, 3., hit the wrong key

OpenStudy (alyssajobug):

3.0625

OpenStudy (phi):

that is t^2

OpenStudy (alyssajobug):

ya, so t is 1.75

OpenStudy (phi):

of course sqr(49)/sqr(16) = 7/4 so you get a nice number for t= 1 ¾ or 1.75 exactly

OpenStudy (alyssajobug):

t=1.75, so -16(1.75)^2+49=0, so 1.75 is my answer?

OpenStudy (phi):

yes. they wanted it rounded to the nearest 100th, so 1.75 is the answer.

OpenStudy (phi):

good job!

OpenStudy (alyssajobug):

OK! GREAT! I think I get the concept now! If I have any questions on others I'll make sure to ask you, THANK YOU SOOOOO MUCH!!!!!!

OpenStudy (alyssajobug):

An outfielder throws a baseball to the player on third base. The height h of the ball in feet is modeled by the function h(t) = -16t2 + 19t + 5, where t is time in seconds. The third baseman catches the ball when it is 4 feet above the ground. To the nearest tenth of a second, how long was the ball in the air before it was caught? \[-16t^2+19t+5=4\]\[-16t^2+19t+1=0\]\[x=\frac{ -19\pm \sqrt{19^{2}-4\times-16\times1} }{ 2\times-16 }\]\[x=\frac{ -19\pm \sqrt{361-64} }{ -32 }\]\[x=\frac{ -19\pm \sqrt{297} }{ -32 }\]\[x=\frac{ -19\pm 17.2336879396}{ -32 }\]\[x=\frac{ -19 + 17.2336879396}{ -32 } = 0.05519725\]\[x=\frac{ -19 - 17.2336879396}{ -32 } = 1.13230275\] @phi did I do it right here? The answers dont seem to work out, they made the equasion = 2 and 2.09749553...

OpenStudy (phi):

inside the square root you have 361 - 4*-16*1 notice that simplifies to 361 - (-64) = 361+64

OpenStudy (alyssajobug):

so the inside is 425 instead then? oh! wow that threw my whole equasion off...

OpenStudy (alyssajobug):

\[x=\frac{ -19+20.6155281281 }{ -32 } = 0.05048525\] \[x=\frac{ -19-20.6155281281 }{ -32 } = 1.23798525\]

OpenStudy (alyssajobug):

@phi

OpenStudy (phi):

The first number should be negative (and obviously not the answer) for the answer, round the second value the nearest tenth of a second

OpenStudy (phi):

notice if we "plug in" t= 1.23798525 into the original equation h(t) = -16t2 + 19t + 5 you will get h= 4 feet which is the height at which the baseball was caught. So that looks good.

OpenStudy (alyssajobug):

Oh! I was kinda confused because I thought it was supposed to make the equasion 0 but it is 4 ft above the ground! DUH whoops, thanks!